We are trying to solve these problems and aren't sure if we are doing them correctly and the formula for doing them. Our first problem was
-24u4v3x5
__________
-18uv5x2
Thank you for any help and please show us the formula you used to get the correct answer. Thank you again.

Found 3 solutions by checkley75, stanbon, jmg:
Answer by checkley75(3666)

THERE IS NO REAL FORMULAS AS SUCH BUT WHAT YOU NEED TO DO HERE IS TO FIND ALL THE COMMON FACTORS IN THE NUMERATOR & THE DENOMINATOR.
A COMMON FACTOR FOR -18 & -24 IS -6. THUS YOU SHOULD DIVIDE BOTH BY -6 & YOU GET
4/3
FACTOR THE U, V, & X THUS:
U4/U=U3
V^3/V^5=1/V^2
X^5/X^2=X^3
NOW WE PUT ALL THESE TERMS BACK TOGETHER THUS:
4U^3X^3/3V^2

Answer by stanbon(48535)

-24u4v3x5
__________
-18uv5x2
===================
Rewrite with related factors:
(-24/-18)(u^4/u)(v^3/v^5)(x^5/x^2)
Reduce each fraction:
=(4/3)(u^3)(1/v^2)(x^3)
=(4/3)(u^3x^3)/(v^2)
=======
Cheers,
Stan H.

Answer by jmg(22)

You ask for a formula, I guess it would be
u^x/u^y = u^x-y
Another words you just subtract the exponents.
To solve the problem:
Simplify -24/-18 by dividing both by 6, to 4/3
u^4/u^1 = u^3 (subtract 4-1)
v^3/v^5 = v^-2 (subtract 3 - 5)
x^5 - x^2 = x^3
When you put it all together you have
(4/3)u^3v^-2x^3
Which can also be written as a fraction:
4u^3x^3
________
3v^2